Find the domain of the function
f(x)= sqrt25-x
Write the function whose graph is the graph of y=(x+2)^2 but is reflected the x-axis
y=____

Answer by stanbon(75887)

Find the domain of the function
f(x)= sqrt(25-x)
25-x has to be >= 0
-----
Solve: 25-x >= 0
x <= 25
x <= 5
============================
Write the function whose graph is the graph of
y=(x+2)^2 but is reflected the x-axis
----
Ans: y = -(x+2)^2
============================
Cheers,
Stan H.
================
y=____